Circular No. 6528

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cb IN NGC 3510 S. D. Van Dyk, Los Angeles, CA; R. A. Sramek, National Radio As-

tronomy Observatory; M. J. Montes and K. W. Weiler, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute and European Space Agency, report the detection of radio emission from the type-II supernova 1996cb on Dec. 21 with the Very Large Array in configuration A: “A radio-flux density of 2.41 ± 0.14 mJy at 8.435 GHz (wavelength 3.6 cm) and 3.00 ± 0.36 mJy at 14.965 GHz (2.0 cm) is measured at a position of α = 11^{h}03^{m}42^{s}.00, δ = +28^{o}54′14′′.2 (equinox 2000.0), within 1′′ of the optical position reported for the supernova (IAUC 6524). A detection of several mJy has also been obtained at 22.46 GHz (wavelength 1.3 cm), but, due to atmospheric phase instabilities, the flux density is being re-evaluated. Radio monitoring is continuing. Since SN 1996cb is relatively nearby at ∼ 8 Mpc, study at other wavelengths is strongly recommended.”

V2116 OPHIUCHI J. Phillips, J. Greenhill and K. M. Hill, University of Tasmania, report:

“Using the Canopus 1-m telescope and CCD photometer, we have monitored V2116 Oph (= GX 1+4) in V, R, I, and Hα at frequent intervals during 1995 and 1996. The Hα flux, measured with a filter of bandwidth 10 nm, reached a peak value for 1996 on or about Aug. 7 and then decreased to a minimum ∼ 3.5 mag fainter on Oct. 1; it remained near this lower level until the end of observations on Oct. 20. This level is consistent with the absence of any line emission on Oct. 16 (cf. IAUC 6496). The onset of decline coincides to within a few days with that in the BATSE hard-x-ray emission (IAUC 6478). V and I flux also peaked on Aug. 7. V decreased by at least 2 mag and was below threshold (V > 20) during October. The decrease in I was ∼ 0.6 mag, indicating that companion emission dominates in this band.”

COMET 118P/SHOEMAKER-LEVY 4 Total visual magnitude estimates: Dec. 1.89 UT, 12.5 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector); 13.94, 12.7 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 29.50, 12.7 (A. Pearce, Woodend, Vic., Australia, 0.20-m reflector).

1996 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green